HIGH RESOLUTION OVERTONE SPECTROSCOPY OF ACETYLENE-WATER VAN DER WAALS COM-PLEXES

\underline{K. DIDRICHE}, T. FOLDES, M. HERMAN, Service de Chimie quantique et Photophysique CP160/09, Faculté des Sciences, Université Libre de Bruxelles (U.L.B.), Av. Roosevelt, 50, B-1050, Bruxelles, Belgium.

The FANTASIO+ set-up^{a} (for ’Fourier trANsform, Tunable, diode and quadrupole mAss spectrometers interfaced to a Supersonic expansIOn’) was upgraded by implementing a pulsed nozzle and a liquid injector. Experiments were performed to test the new capacities of this improved set-up. The first results will be presented, among which the CW-cavity ring down spectra of the ^{12}C_{2}H_{2}-H_{2}O and ^{12}C_{2}H_{2}-D_{2}O absorption bands in the 2CH excitation range (1.5 μm). The first analyses of these spectra will be presented.

aDidriche et al., Mol. Phys. 2010, 108, 2158-2164